Find three consecutive integers that sun to -42
Answer by ikleyn(52786)

.

            Watch my steps . . .


(x-1) + x + (x+1) = -42


3x                = -42


 x                = -42/3 = -14     the middle integer.


ANSWER.  The three consecutive integers  are -15, -14 and -13.


CHECK.   Make the check on your own.


Notice. My solution was so simple, because I selected the MIDDLE number of three consecutive integers as my major unknown.